A coffee merchant has coffee beans that sell for $9.00 per pound and $12.00 per pound. The two types are to be mixed to create 100 pounds of a mixture that will sell for $11.25 pound. How much of each type of bean should be used in the mixture?
Answer by stanbon(75887)

A coffee merchant has coffee beans that sell for $9.00 per pound and $12.00 per pound. The two types are to be mixed to create 100 pounds of a mixture that will sell for $11.25 pound. How much of each type of bean should be used in the mixture?
--------
Let the amount of $9.00 coffee be "x"lbs.
The value of this coffee is 9x dollars.
The amount of $12 coffee is "100-x" lbs.
The value of this coffee is 12(100-x)=1200-12x dollars
The amount of mixed coffee is 100 lbs.
The value of the mixed is 11.25(100)= $1125
EQUATION:
value + value = value
9x + 1200-12x=1125
-3x=-75
x=25 lbs. (amount of $9 coffee)
100-x=75 lbs. (amount of $12 coffee
Cheers,
Stan H.